4. The grades for a large precalculus class are normally distributed with mean 60 and standard deviation 10.

(a) What should the A/B line be if only 5% of the class is to get A’s?

(b) If the D/C line is 65, what should the C/B line be if 25% of the class is to get C’s?

5. Of Maryland citizens, 85% hate country music. 400 citizens are selected at random.

(a) What is the expected number of people who hate country music?

(b) Approximate the probability that more than 350 hate country music.

6. The number of soda cans bought per week by a person is distributed with mean 3 and standard deviation 1.7. We observe the the number of cans 100 particular people buy during one particular week.

What is the probability that the average number of cans they buy is between 2.9 and 3.2?